An alpha particle can be produced in certain radioactive decays of nuclei and consists of two protons and two neutrons. The particle has a charge of q = +2e and a mass of 4.00 u, where u is the atomic mass unit, with 1 u = 1.661 ✕ 10−27 kg. Suppose an alpha particle travels in a circular path of radius 4.60 cm in a uniform magnetic field with B= 1.70 T. Calculate the following values.

(a) its speed
  m/s

(b) its period of revolution
s

(c) its kinetic energy
eV

(d) the potential difference through which it would have to be accelerated to achieve this energy
V
Answer #1

a)

r = 0.046 m

B = 1.70 T

m = mass = 4 u = 4 x 1.661 x 10-27 kg

q = charge = 2 e = 2 x 1.6 x 10-19

v = speed

using the formula

r = mv /qB

v = qBr/m

v = (2 x 1.6 x 10-19) (1.70) (0.046) /(4 x 1.661 x 10-27)

v = 3.8 x 106 m/s

b)

period of revolution is given as

T = 2\pim/(qB)

T = 2 (3.14) (4 x 1.661 x 10-27 )/((2 x 1.6 x 10-19) (1.70)) = 7.7 x 10-8 sec

c)

kinetic energy is given as

KE = (0.5) m v2 = (0.5) (4 x 1.661 x 10-27 ) (3.8 x 106 )2 = 4.8 x 10-14 J = 4.8 x 10-14/(1.6 x 10-19) = 3 x 105 eV

d)

using conservation of energy

electric potential energy = kinetic energy

q\DeltaV = KE

(2 x 1.6 x 10-19) \DeltaV = 4.8 x 10-14

\DeltaV = 1.5 x 105 volts